# Large Language Model-informed ECG Dual Attention Network for Heart Failure Risk Prediction

### 15 Mar 2024 v1

@@ Abstract
Heart failure (HF) poses a significant public health challenge due to its rising global mortality rate. Addressing this issue through early diagnosis and prevention could significantly reduce the disease’s impact. This work introduces a methodology for HF risk prediction using clinically acquired 12-lead electrocardiograms (ECGs). We present a novel, lightweight dual-attention ECG network designed to capture complex ECG features essential for early HF prediction, despite the notable imbalance between low and high-risk groups. The network features a cross-lead attention module and twelve lead-specific temporal attention modules to capture cross-lead interactions and local temporal dynamics within each lead. To prevent model overfitting from limited training data, we leverage a large language model (LLM) with a public ECG-Report dataset for pretraining on an ECG-report alignment task. The network is then fine-tuned for HF risk prediction using two specific cohorts from the UK Biobank study, focusing on patients with hypertension (UKB-HYP) and those who have had a myocardial infarction (UKB-MI). Our findings show that LLM-informed pretraining significantly improves the network’s HF risk prediction capability in these cohorts. Moreover, the dual-attention mechanism enhances interpretability and predictive performance, ensuring a transparent and reliable prediction process. The method outperforms existing models, achieving average C-index scores of 0.6349 and 0.5805 on the UKB-HYP and UKB-MI test sets, respectively. This performance demonstrates our approach’s effectiveness in managing complex clinical ECG data and its potential to improve HF risk assessment across various populations.
@@
Preventing HF early is crucial to reduce its health and economic impacts, yet HF diagnosis often occurs late when patients have already developed serious symptoms [4, 5]. risk [6, 7, 8, 9, 10]. A promising strategy for improving heart failure (HF) management is the development of risk prediction models for future HF events. These models generate a risk score for patients over a specific timeframe, taking into consideration of the patient’s specific characteristics. With such a personalized risk assessment, more tailored HF management strategies and/or treatment recommendations can be provided. In this context, the low-cost 12-lead electrocardiogram (ECG), a medical test commonly used in clinical practice, serves as a valuable resource for evaluating a patient’s cardiovascular health and uncovering the risk. Recent studies have already found that several markers detected from clinically acquired 12‐lead ECG are associated with future HF events, such as prolonged QRS duration [6, 7, 8, 9, 10], conduction disorders (left, right bundle-branch blocks) [11, 12]. A significant limitation of these previous research lies in its reliance on a limited set of biomarkers (e.g., QRS duration), which are identified through predetermined rules based on ECG data. Moreover, much of this research has employed simple linear models for modeling the risk ratio associated with HF. While linear models offer a straightforward and interpretable framework for risk assessment, they may fail to capture the intricate, complex subtleties and nuances embedded within the ECG for the early-stage risk prediction of HF.
@@ II-A Risk prediction
Risk prediction aims to estimate the chance that a patient will have a certain event in an infinitesimal time interval : , given that the event has not occurred before. A common approach is the Cox proportional hazards formulation (CoxPH) [17]. The CoxPH model is a statistical model for predicting the time-to-event outcomes based on the assumption that the hazard function is proportional to a set of features or variables associated with the subject being studied. Mathematically, it can be defined as: , where models the hazard rate at which events occur at a time taking the subject information into account, is the baseline hazard function shared by all observations, depending only on the time , and is the risk score for observation . Conventional CoxPH models assume that the input observation is a set of covariates (e.g., sex, age, medical history, smoking) and constrain the function to the linear form: with being the weights to the set of covariates in [17]. In the existing body of literature pertaining to heart failure risk prediction, the predominant methodologies employ a predefined set of variables to construct a linear model, subsequently isolating variables that exhibit a high correlation for the purpose of risk prediction. This approach includes parameters such as prolonged QRS duration [6, 7, 8, 9, 10], various conduction disorders (specifically left and right bundle-branch blocks) [11, 12], elongated QT intervals [7], abnormalities in QRS/T angles and T wave patterns [18, 19, 20, 21, 7], and ST-segment depression in the V5 precordial lead [7, 22].
@@ III-B1 Lead attention
where is the attention weight matrix and is the feature dimension (). We use the same input for . In practice, for computational efficiency, following [37], we apply the multi-head attention trick, which first linearly projects the queries, keys, and values times with different learnable linear projection matrices to lower dimensions () respectively to compute the attention matrix and re-weight the projected value matrix, then concatenate all the output heads to recover the feature dimension.
@@
where and are weights and biases of two linear layers in FFN, LayerNorm is a layer normalization layer [38] for feature normalization. In this way, the output feature is adjusted with all the information from other leads into account.
III-B2 12-lead temporal attention
The structure of the temporal attention module is very similar to the lead attention. The only difference is that now we have separate attention modules for different lead features so that each module is adapted to a specific lead to extract temporal dynamic features locally. Specifically, we split the feature h along the lead dimension, and for each lead feature , we first add sinusoid positional encoding information along the time domain and then process each lead with a separate temporal attention module where each time point is treated as a token. In other words, the input becomes a sequence of -dimensional features with a sequence length of and the corresponding temporal attention matrix for lead becomes: . We also send the reweighted output to an FFN module for feature refinement for each lead, similar to the lead attention module process. The output of the temporal attention is a concatenation of output from the 12 TA modules:
@@ III-C1 Large language model informed model pre-training
We employ an LLM to extract the knowledge embedded in ECG reports. Specifically, we use the medical domain language model: BioClinical BERT [42] 111https://huggingface.co/emilyalsentzer/Bio_ClinicalBERT, which has been trained on a large number of electronic health records from MIMIC III [43]. We consider two ways of extracting text embeddings :
Latent text code from raw ECG report (raw): For a piece of ECG report (in English) 444Since the original reports are written in a mixture of German and English, we used the open-source machine translation tool: EasyNMT 222https://github.com/UKPLab/EasyNMT for batch translation following [34] and then used ChatGPT333https://chat.openai.com/ to further refine those failed cases with a prompt: translate the ECG report into English:., we simply feed it to the LLM to get , following [35, 32].
Latent text code from structured ECG report weighted by confidence (structured with confidence): The original PTB-XL dataset also provides ECG-SCP codes generated from ECG reports. Specifically, each report has been abstracted to a set of SCP codes with SCP-ECG statement description and confidence score (annotated by human experts). In this case, we build a structured sentence with linked SCP statement category information and SCP description for each SCP code: = ‘Statement Category(SCP)}:SCP-ECG Statement Description(SCP)}’555Database containing SCP code-statement mappings can be found at https://physionet.org/content/ptb-xl/1.0.1/scp_statements.csv and https://physionet.org/content/ptb-xl/1.0.1/ptbxl_database.csv. For example, as shown in Fig 2(a), given an SCP code: LNGQT, the input becomes: ‘other ST-T descriptive statements: long QT-interval’. We send this type of structured input to the LLM, which gives an embedding for each SCP code. It is important to note that a single recording may encompass multiple SCP codes. In that case, multiple SCP embeddings are initially derived. These embeddings are then aggregated with weights based on corresponding confidence scores to get the text embedding for the corresponding ECG: , where is the sum of all confidence scores for all corresponding statements for the input ECG signal.
@@ V-C Comparison study between traditional ECG parameter-based risk prediction vs ECG dual attention
We further compare our method to the traditional method with well-established ECG parameters. Specifically, we collect a set of ECG parameters from the ECG signals, which have been identified in previous studies for incident HF prediction and mortality estimation [6, 8, 9, 11, 12, 18, 19, 20, 21, 7, 22]. All of these parameters were automatically extracted by the ECG device (using the GE CardioSoft V6 666https://biobank.ndph.ox.ac.uk/ukb/ukb/docs/CardiosoftFormatECG.pdf) with supporting evidence from previous work: 1) Ventricular rate; 2) Left-axis deviation; 3) Right-axis deviation; 4) Prolonged P-wave duration () [22]; 5) Prolonged PR interval () [22]; 6) Prolonged QRS duration () [22]; 7) Prolonged QT interval ( using the Framingham formula) [52]; 8) Delayed intrinsicoid deflection (DID time) (the maximum value in leads V5 and V6 ) [53]; 9) Abnormal P-wave axis (values outside the range of and ) [54]; 10) Left ventricular hypertrophy [55]; 11) Abnormal QRS-T angle () [22]; 12) Low QRS voltage [56]; 13) ST-T abnormality [56]; 14) Right bundle‐branch block [56]; 15) Left bundle‐branch block [56].
@@ VI Discussion
For better understanding, we visualizes the cross-lead and temporal activation maps of two high-risk cases in Figure 6. It can be observed that cross-lead module identify the uniqueness or exploit synergy among various leads, whereas temporal lead attention focuses on identifying local areas in each lead important to the prediction. For example, Leads III and II stand out in the cross-lead attention maps (a) and (b) respectively, signaling unique pathological pattern (prolonged negative QRS in lead III, see (d) and abnormal noises in lead II, see (e), in contrast to other leads. Additionally, the orange diagonal clusters in panel (b) showcases good R wave progression from V2 to V6, see (e). The temporal attention maps in (d) and (e) reveal more intricate details, showing the neural network’s capability to focus on clinically significant features such as P, Q, S, T waves, and R peaks, even amidst noise, and highlight abnormalities like abnormal R peaks, T wave iregularities and prolonged QRS and QT intervals. This demonstrates that the deep learning model is adept at autonomously discovering clinically relevant metrics from ECG data without explicit instruction during its training phase. Future research will aim at collaborating with medical professionals to unearth potential novel biomarkers using these visual tools.
@@ VII Conclusion
This paper presents a study with a novel ECG dual attention network for ECG-based heart failure risk prediction. This network distinguishes itself through its unique blend of being both lightweight and efficient, outperforming existing models in the field. Its standout feature is the ability to generate not just a quantitative risk score but also to provide a qualitative interpretation of its internal processes. This is achieved through the generation of attention visualization maps, which span across and delve within individual leads, offering a granular view of the network focus and decision-making process. We hope to contribute to the development of more transparent and interpretable AI-assisted systems, fostering trust and facilitating broader adoption in clinical settings. Additionally, the study underscores the importance of pretraining in addressing the challenges posed by limited data in risk prediction models. We suggest a multi-modal pretraining approach that leverages external public ECG reports and confidence scores from diverse populations, combined with advanced large language models, to enhance the model’s data efficiency and effectiveness.

### 22 Mar 2024 v2

@@ Abstract
Heart failure (HF) poses a significant public health challenge, with a rising global mortality rate. Early detection and prevention of HF could significantly reduce its impact. We introduce a novel methodology for predicting HF risk using 12-lead electrocardiograms (ECGs). We present a novel, lightweight dual-attention ECG network designed to capture complex ECG features essential for early HF risk prediction, despite the notable imbalance between low and high-risk groups. This network incorporates a cross-lead attention module and twelve lead-specific temporal attention modules, focusing on cross-lead interactions and each lead’s local dynamics. To further alleviate model overfitting, we leverage a large language model (LLM) with a public ECG-Report dataset for pretraining on an ECG-report alignment task. The network is then fine-tuned for HF risk prediction using two specific cohorts from the UK Biobank study, focusing on patients with hypertension (UKB-HYP) and those who have had a myocardial infarction (UKB-MI).The results reveal that LLM-informed pre-training substantially enhances HF risk prediction in these cohorts. The dual-attention design not only improves interpretability but also predictive accuracy, outperforming existing competitive methods with C-index scores of 0.6349 for UKB-HYP and 0.5805 for UKB-MI. This demonstrates our method’s potential in advancing HF risk assessment with clinical complex ECG data.
@@
Preventing HF early is crucial to reduce its health and economic impacts, yet HF diagnosis often occurs late when patients have already developed serious symptoms [4, 5]. risk [6, 7, 8, 9, 10]. A promising strategy for improving HF management is the development of risk prediction models for future HF events. These models generate a risk score for patients over a specific timeframe, taking into consideration of the patient’s specific characteristics. With such a personalized risk assessment, more tailored HF management strategies and/or treatment recommendations can be provided. In this context, the low-cost 12-lead electrocardiogram (ECG), a medical test commonly used in clinical practice, serves as a valuable resource for evaluating a patient’s cardiovascular health and uncovering the risk. Recent studies have already found that several markers detected from clinically acquired 12‐lead ECG are associated with future HF events, such as prolonged QRS duration [6, 7, 8, 9, 10], conduction disorders (left, right bundle-branch blocks) [11, 12]. A significant limitation of these previous research lies in its reliance on a limited set of biomarkers (e.g., QRS duration), which are identified through predetermined rules based on ECG data. Moreover, much of this research has employed simple linear models for modeling the risk ratio associated with HF. While linear models offer a straightforward and interpretable framework for risk assessment, they may fail to capture the intricate, complex subtleties and nuances embedded within the ECG for the early-stage risk prediction of HF.
@@ II-A Risk prediction
Risk prediction aims to estimate the chance that a patient will have a certain event in an infinitesimal time interval : , given that the event has not occurred before. A common approach is the Cox proportional hazards formulation (CoxPH) [17]. The CoxPH model is a statistical model for predicting the time-to-event outcomes based on the assumption that the hazard function is proportional to a set of features or variables associated with the subject being studied. Mathematically, it can be defined as: , where models the hazard rate at which events occur at a time taking the subject information into account, is the baseline hazard function shared by all observations, depending only on the time , and is the risk score for observation . Conventional CoxPH models assume that the input observation is a set of covariates (e.g., sex, age, medical history, smoking) and constrain the function to the linear form: with being the weights to the set of covariates in [17]. In the existing body of literature pertaining to HF risk prediction, the predominant methodologies employ a predefined set of variables to construct a linear model, subsequently isolating variables that exhibit a high correlation for the purpose of risk prediction. This approach includes parameters such as prolonged QRS duration [6, 7, 8, 9, 10], various conduction disorders (specifically left and right bundle-branch blocks) [11, 12], elongated QT intervals [7], abnormalities in QRS/T angles and T wave patterns [18, 19, 20, 21, 7], and ST-segment depression in the V5 precordial lead [7, 22].
@@ III-B1 Lead attention
where is the attention weight matrix and is the feature dimension (). We use the same input for computing 111In practice, for computational efficiency, following [37], we apply the multi-head attention trick, which first linearly projects the queries, keys, and values times with different learnable linear projection matrices to lower dimensions () respectively to compute the attention matrix and re-weight the projected value matrix, then concatenate all the output heads to recover the feature dimension..
@@
where and are weights and biases of two linear layers in FFN; LayerNorm is a layer normalization layer [38] for feature normalization. In this way, the output feature is adjusted with all the information from other leads into account.
III-B2 12-lead temporal attention (TA)
The structure of the temporal attention module is very similar to the lead attention. The only difference is that now we have separate attention modules for different lead features so that each module is adapted to a specific lead to extract temporal dynamic features locally. Specifically, we split the feature h along the lead dimension, and for each lead feature , we first add sinusoid positional encoding information along the time domain and then process each lead with a separate temporal attention module where each time point is treated as a token. In other words, the input becomes a sequence of -dimensional features with a sequence length of and the corresponding temporal attention matrix for lead becomes: . Similar to the lead attention module process, we reweight and normalize the lead feature with the generated temporal attention matrix and then employ an FFN module for feature refinement for each lead. The output of the temporal attention is a concatenation of output from the 12-lead temporal attention modules. The whole process can be defined as follows:
@@ III-C1 Large language model informed model pre-training
We employ an LLM to extract the knowledge embedded in ECG reports. Specifically, we use the medical domain language model: BioClinical BERT [42] 222https://huggingface.co/emilyalsentzer/Bio_ClinicalBERT, which has been trained on a large number of electronic health records from MIMIC III [43]. We consider two ways of extracting text embeddings :
Latent text code from raw ECG report (raw): For a piece of ECG report (in English) 555Since the original reports are written in a mixture of German and English, we used the open-source machine translation tool: EasyNMT 333https://github.com/UKPLab/EasyNMT for batch translation following [34] and then used ChatGPT444https://chat.openai.com/ to further refine those failed cases with a prompt: translate the ECG report into English:., we simply feed it to the LLM to get , following [35, 32].
Latent text code from structured ECG report weighted by confidence (structured with confidence): The original PTB-XL dataset also provides ECG-SCP codes generated from ECG reports. Specifically, each report has been abstracted to a set of SCP codes with SCP-ECG statement description and confidence score (annotated by human experts). In this case, we build a structured sentence with linked SCP statement category information and SCP description for each SCP code: = ‘Statement Category(SCP)}:SCP-ECG Statement Description(SCP)}’666Database containing SCP code-statement mappings can be found at https://physionet.org/content/ptb-xl/1.0.1/scp_statements.csv and https://physionet.org/content/ptb-xl/1.0.1/ptbxl_database.csv. For example, as shown in Fig 2(a), given an SCP code: LNGQT, the input becomes: ‘other ST-T descriptive statements: long QT-interval’. We send this type of structured input to the LLM, which gives an embedding for each SCP code. It is important to note that a single recording may encompass multiple SCP codes. In that case, multiple SCP embeddings are initially derived. These embeddings are then aggregated with weights based on corresponding confidence scores to get the text embedding for the corresponding ECG: , where is the sum of all confidence scores for all corresponding statements for the input ECG signal.
@@ V-C Comparison study between traditional ECG parameter-based risk prediction vs ECG dual attention
We further compare our method to the traditional method with well-established ECG parameters. Specifically, we collect a set of ECG parameters from the ECG signals, which have been identified in previous studies for incident HF prediction and mortality estimation [6, 8, 9, 11, 12, 18, 19, 20, 21, 7, 22]. All of these parameters were automatically extracted by the ECG device (using the GE CardioSoft V6 777https://biobank.ndph.ox.ac.uk/ukb/ukb/docs/CardiosoftFormatECG.pdf) with supporting evidence from previous work: 1) Ventricular rate; 2) Left-axis deviation; 3) Right-axis deviation; 4) Prolonged P-wave duration () [22]; 5) Prolonged PR interval () [22]; 6) Prolonged QRS duration () [22]; 7) Prolonged QT interval ( using the Framingham formula) [52]; 8) Delayed intrinsicoid deflection (DID time) (the maximum value in leads V5 and V6 ) [53]; 9) Abnormal P-wave axis (values outside the range of and ) [54]; 10) Left ventricular hypertrophy [55]; 11) Abnormal QRS-T angle () [22]; 12) Low QRS voltage [56]; 13) ST-T abnormality [56]; 14) Right bundle‐branch block [56]; 15) Left bundle‐branch block [56].
@@ VI Discussion
For better understanding, we visualizes the cross-lead and temporal activation maps of two high-risk cases in Figure 6. For temporal ones, we apply GradCAM [61] to mapping the low-resolution temporal attention maps back to the original input signal level (see appendix for more details). It can be observed that cross-lead module provides an overview to identify the uniqueness or exploit synergy or cross-lead interactions among various leads, whereas temporal lead attention focuses on identifying local areas in each lead important to the prediction. For example, Leads III and II stand out in the cross-lead attention maps (a) and (b) respectively, highlighting unique pathological pattern (prolonged negative QRS in lead III, see (d) or strong abnormal noises in lead II, see original signals in (e), in contrast to other leads. Additionally, the orange diagonal clusters in panel (b) showcases good R wave progression from V2 to V6, see (e). On the other hand, the temporal attention maps in (d) and (e) reveal more intricate details, showing the neural network’s capability to focus on clinically significant features such as P, Q, S, T waves, and R peaks, even with strong noise presented, and highlight pathological abnormalities like abnormal R peaks, T wave iregularities and prolonged QRS and QT intervals. This implies that the dual attention network can autonomously discover clinically relevant biomarkers from ECG data without explicit being taught during its training phase. Future research will aim at collaborating with medical professionals to verify potential novel biomarkers using these visual tools.
@@ VII Conclusion
This paper presents a study with a novel ECG dual attention network for ECG-based HF risk prediction. This network distinguishes itself through its unique blend of being both lightweight and efficient, outperforming existing models in the field. Its standout feature is the ability to generate not just a quantitative risk score but also to provide a qualitative interpretation of its internal processes. This is achieved through the generation of attention visualization maps, which span across and delve within individual leads, offering a granular view of the network focus and decision-making process. We hope to contribute to the development of more transparent and interpretable AI-assisted systems, fostering trust and facilitating broader adoption in clinical settings. Additionally, the study presents a multi-modal pretraining approach for risk prediction models, which leverages external public ECG reports and confidence scores from diverse populations, combined with advanced large language models, to address the challenges posed by limited future event data in risk prediction tasks.